# COOCO - Common Objects Out-of-Context - Semantic Violation in Scenes: Investigating Multimodal Context in Referential Communication

### v1

@@ Abstract
Natural scenes provide us with rich contexts for object recognition and reference. In particular, knowing what type of scene one is looking at generates expectations about which objects will occur, and what their spatial configuration should be. Do Vision-Language Models (VLMs) learn to rely on scene contexts in a similar way, when generating references to objects? To address this question, we introduce the Common Objects Out-of-Context (COOCO) dataset and test to what extent VLMs rely on scene context to refer to objects under different degrees of scene-object congruency, and different perturbations. Our findings show that models leverage scene context adaptively, depending on both the semantic relatedness between object and scene and the level of noise. In particular, models rely more on context under high target-scene congruence or when objects are degraded. Attention analysis reveals that successful object categorisation involves increased focus on the target in mid-level layers, especially under moderate noise, suggesting that VLMs dynamically balance local and contextual information for reference generation. We make our dataset, code and models available at https://github.com/cs-nlp-uu/scenereg.
@@ 1 Introduction
Objects rarely appear in isolation. Over time, humans build expectations of seeing certain objects in certain contexts. For instance, we expect to see a laptop on a desk in an office, but not a large piece of ham (Figure 1). Studies of human perception in visual scenes Võ (2021); Torralba et al. (2006); Coco et al. (2016) suggest that scenes exhibit both semantic and syntactic regularities Biederman et al. (1982); Võ (2021). Violations of either or both of these result in processing difficulties  Ganis and Kutas (2003); Öhlschläger and Võ (2016); Võ (2021).
@@
Figure 1: An original image of an office scene from the COCO dataset and modified version from our dataset, in which the laptop is replaced with an object that has low relevance to the scene. Bottom: versions of the modified image with Gaussian noise mask (level = 0.5) on the context and on the target.
In this work, we are interested in how Vision-Language Models (VLMs) behave in the presence of semantic violations in visual scenes when referring to objects. Such an analysis offers insights into the extent to which models learn to leverage contextual visual information. We focus on object naming, that is, the task of determining the category of an object, a fundamental part of Referring Expression Generation (REG; Krahmer and Van Deemter, 2012). The literature on scene perception suggests that scene semantics mediates human viewers’ expectations about the objects in a scene, and viewers deploy their attentional resources accordingly. Scene semantics also influence how humans describe objects in scenes Hwang et al. (2011); de Groot et al. (2017, 2015). Do VLMs similarly utilise scene-level visual information when generating descriptions of objects? To study this, we compare the naming of objects in semantically (in)congruent scenes. If scene-level information guides model choices, then there should be systematic differences between these two conditions in the way models refer to target referents. Previous work has shown the facilitating effects of context in REG, even in the case of perturbations made on images (Junker and Zarrieß, 2024). However, this work does not control for the impact of semantic (in)congruity, and relies on small-scale, purposely-trained models or on artificial images (Junker and Zarrieß, 2025).
We develop a new dataset, Common Objects Out-of-Context (COOCO), which manipulates natural images to introduce objects of varying degrees of consistency with the scene (Figure 1). Similarly to Junker and Zarrieß (2024), we further manipulate the visibility of the target referent by introducing different degrees of noise. We summarise the contributions of our study as follows.
1. We present COOCO, a novel dataset to evaluate multimodal models’ capabilities for integrating object-level and scene-level visual information, under both congruent and incongruent conditions.
2. We investigate the impact of scene context on object identification in state-of-the-art VLMs. We show that under conditions where target objects violate scene semantics, scene context acts as a distractor, while it has a clear facilitating effect when the target is congruent with the scene. This also makes models resilient to noise when it is applied to the target region.
3. We investigate how attention is distributed by VLMs across scene components under different conditions of visual semantic congruency, when generating referring expressions.
@@ 2 Related Work
Context in Referring Expression Generation.
@@
Referring Expression Generation (REG) is the task of generating identifying descriptions for a target object. Early REG approaches relying on handcrafted algorithms Krahmer and Van Deemter (2012) incorporated contextual influences by considering the properties that distinguish a target referent from its distractors, with some research also incorporating constraints to guide coherent categorisation for references to plurals, and the use of landmarks for spatial expressions Schüz et al. (2023). More recent neural models for visual REG learn mappings from raw perceptual inputs to linguistic expressions. Here too there has been limited attention on the ability of models to leverage global scene information Schüz et al. (2023). This gap has motivated recent investigations into how scene information may enhance robustness in challenging settings Schüz and Zarrieß (2023); Junker and Zarrieß (2025). In particular, Junker and Zarrieß (2024) show that models rely on scene context to generate references even when the target object is fully occluded, indicating that scene context serves not only as a disambiguating constraint but also as a supportive resource for robust reference generation.
Context in Human Object Recognition.
@@
Human object recognition is strongly influenced by contextual expectations Bar (2004); Oliva and Torralba (2007); Võ (2021); Peelen et al. (2023), including both global scene properties (e.g., gist, layout) and local cues (e.g., anchor objects, co-occurring items) Lauer et al. (2018, 2021); Boettcher et al. (2018). Evidence that visual search is semantically constrained comes from viewing tasks, where viewers fixate more on semantically informative (but not necessarily visually salient) regions, and semantic similarity is a strong predictor of gaze trajectories. Both highly expected and highly incongruent objects attract more attention than moderately fitting ones Damiano et al. (2024).
Datasets.
@@
Widely used vision and language datasets for referring expression and image caption generation, such as COCO Lin et al. (2015), RefCOCO Kazemzadeh et al. (2014); Yu et al. (2016a) and VisualGenome Krishna et al. (2016), lack annotations for object-scene relatedness and do not enable research on semantic violations in scenes. A number of scene violation datasets have been created within the vision community. Prominent examples include ObjAct Shir et al. (2021), Event Task Dresang et al. (2019), Out-of-Context Bomatter et al. (2021), Cut-and-paste dataset Yun et al. (2021), and SCEGRAM Öhlschläger and Võ (2016). However, these datasets tend to be relatively small-scale and do not incorporate degrees of semantic violation or occlusion. In this work, we present a new, large-scale dataset for semantic violations in visual scenes, building on existing resources such as COCO-Search18 Chen et al. (2021), THINGSPlus Hebart et al. (2019); Stoinski et al. (2023) and SUN Xiao et al. (2010).
@@ 3 Dataset Creation
We introduce the Common Objects Out-of-Context (COOCO) dataset. We design this dataset to provide a collection of images where the semantic coherence of a scene is disrupted by the presence of a target object with low, medium or high semantic relatedness to the scene type. COOCO builds upon a framework previously established in the visual perception literature with the SCEGRAM dataset Öhlschläger and Võ (2017), scaling it up to meet the needs of NLP and computer vision research through a novel methodology.
@@
We begin with the publicly available subset of COCO-Search18 Chen et al. (2021), which includes bounding boxes for target objects. This dataset excludes images with people or animals, categories which are known to significantly influence visual attention  Clarke et al. (2013); End and Gamer (2017). Excluding such cues, therefore, avoids confounding factors and enables us to study object-scene semantic relationships. COCO-Search18 further excludes images with targets with multiple object instances, controlling for target size, location, and image aspect ratio. We retained the 2,241 images where the target object is present.
@@
For each image, we predict the scene label using a Vision Transformer (ViT) [^1] https://huggingface.co/google/vit-base-patch16-224 Dosovitskiy et al. (2021) finetuned on SUN-397 Xiao et al. (2010). We then reduce the SUN label set (397 labels) to 25 (listed in Appendix A), selected to balance frequency and semantic diversity and reclassified the images with the same model, excluding all but the selected labels. This approach mitigates risks of category under-representation and excessive variance. This step provided a scene label for each image.
@@
We developed an inpainting pipeline to create versions of images where the target object is replaced with alternatives of varying congruence with the scene (low and medium). We also retain the original image, and additionally include controls where (a) the target is replaced with an object has high relatedness to the scene, and (b) the target is replaced with an object of the same category as the original. The latter are included to verify if the effects found in the experiments are due to artefacts caused by inpainting in general, or due to the introduction of incongruent object types. Figure 2 shows an example of a scene and its corresponding variants in COOCO.
New objects are determined using human-generated norms in THINGSplus Stoinski et al. (2023), which includes typicality ratings (the degree of representativeness of an object for higher-level categories) and ratings of the real-world size of objects. We retain only typical objects (typicality ranging from 0.3 to 1) and remove animate objects to ensure consistency with COCO-Search18. We then match COCO-Search18 targets to candidate object types in THINGSPlus with similar real-world size norms. Candidates differing in size from the targets by more than 25 units are excluded.
Our approach to computing semantic relatedness between candidate replacements and scene labels follows Hayes and Henderson (2021), who found strong correlations between embeddings in ConceptNet Numberbatch Speer et al. (2018) and human visual attention shifts. We use the cosine similarity between scene and object labels to filter low-, medium- and high-similarity candidates.
@@
Given an image, a scene label and a target object with its bounding region and segmentation-based mask obtained from COCO-Search18 we remove the target from the image, and generate 15 new versions of the scene where the target is replaced with objects having low, medium or high-similarity to the scene label (see Figure 2). We also generate an image containing the target as in the original image as a control condition. Only images that pass a verification test are retained. Full details of the inpainting pipeline and verification tests are in Appendix A.
Final dataset
The final COOCO dataset comprises 18,395 images: 1,862 original images with corresponding object-removed versions, 5,480 generated images with low relatedness, 5,572 with medium relatedness, 1,771 with high relatedness, and 1,848 images for the same-target condition.
@@ 4 Experimental Setup
We design our experiments to address key research questions concerning how VLMs leverage contextual information for object naming, an essential component of Referring Expression Generation. In particular, we evaluate the extent to which these models rely on scene context when faced with varying degrees of semantic congruence between an object and its background; and on how robustly they perform REG under different levels and configurations of visual noise. With respect to the second of these questions, we address the findings of Junker and Zarrieß (2024) more systematically, using a carefully designed dataset and testing a range of state-of-the-art pretrained VLMs.
To achieve this, each model is provided with an image and a specified target region, and is prompted to describe the object within that area using a structured text prompt.
@@ 4.1 Models
Molmo Deitke et al. (2024) (7B) is trained on a diverse dataset including ca. 2.3M human-annotated pairs of questions and references to image regions based on 2D points, from 428k images.
@@
LLaVA-OneVision Li et al. (2025) (0.5B and 7B) was trained on data with bounding boxes, incorporating RefCOCO Kazemzadeh et al. (2014); Yu et al. (2016b)[>50k samples] and Visual Genome Krishna et al. (2016)[86k] for single-image settings, and multi-image datasets including WebQA Chang et al. (2022)[9.3k]. The model is instruction-tuned to accept spatial information directly in text input in the form of normalised bounding box coordinates.
@@
For detailed descriptions of preprocessing steps and region-of-interest (ROI) prompt formats for each model, refer to Appendix B.
Prompts.
We test models with structured prompts (details in Appendix B), which include an explicit reference to a region, for example:
What is the object in this part of the image [$x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]? Answer with the object’s name only. No extra text.
4.2 Noise injection.
We add Gaussian noise ($\lambda=0.0,0.5,1.0$), ranging from mild distortion to complete information loss, to the target region, the context region or the entire image (‘All noise’). Pixel values in the affected area are perturbed with noise sampled from a normal distribution and clipped to stay within valid image ranges.
@@ 4.3 Metrics
RefCLIPScore Hessel et al. (2022) is the harmonic mean between (1) the CLIPScore Hessel et al. (2022), which captures the semantic alignment between the generated caption and the image; and (2) the highest cosine similarity between the CLIP Radford et al. (2021) embeddings of the generated caption and any reference caption.
Text-Based Semantic Similarity, in terms of cosine similarity between CLIP text embeddings, which directly assesses the semantic closeness between the model’s output and the target referent label, as well as between the model’s output and the scene label.
Accuracy, defined as 1 if the cosine similarity between CLIP text embeddings of the model’s output and the target label exceeds a threshold (0.9), and 0 otherwise. [^2] We also computed a hard accuracy metric based on Levenshtein similarity, which quantifies the character-level edit distance between generated and reference captions. Since results are very similar to those obtained with the cosine-based metric, we do not report them here.
5 Results
In this section, we report and discuss the main results across all models and experimental conditions in COOCO. We subsequently turn (Sections 6 and 7) to an in-depth exploration of LLaVA-OneVision, which is one of the best-performing models.
Figure 3: RefCLIPScores per model at noise level 0 across relatedness conditions. Models perform best in the original and same target conditions, with performance declining as target relatedness decreases. Warm-colored models show irregular trends, while cool-colored models exhibit a consistent drop.

### v2

# Common Objects Out of Context (COOCo): Investigating Multimodal Context and Semantic Scene Violations in Referential Communication

@@ Abstract
To what degree and under what conditions do VLMs rely on scene context when generating references to objects? To address this question, we introduce the Common Objects Out-of-Context (COOCo) dataset and conduct experiments on several VLMs under different degrees of scene–object congruency and noise. We find that models leverage scene context adaptively, depending on scene-object semantic relatedness and noise level. Based on these consistent trends across models, we turn to the question of how VLM attention patterns change as a function of target-scene semantic fit, and to what degree these patterns are predictive of categorisation accuracy. We find that successful object categorisation is associated with increased mid-layer attention to the target. We also find a non-monotonic dependency on semantic fit, with attention dropping at moderate fit and increasing for both low and high fit. These results suggest that VLMs dynamically balance local and contextual information for reference generation.
Dataset and code are available here: https://github.com/cs-nlp-uu/scenereg.
@@ 1 Introduction
Objects rarely appear in isolation. Over time, humans build expectations of seeing certain objects in certain contexts. For instance, we expect to see a laptop on a desk in an office, but not a large piece of ham (Figure 1). Studies of human scene perception Võ (2021); Torralba et al. (2006); Coco et al. (2016) suggest that scenes exhibit both semantic and syntactic regularities Biederman et al. (1982); Võ (2021). Violations of either or both of these result in processing difficulties  Ganis and Kutas (2003); Öhlschläger and Võ (2016); Võ (2021).
In this work, we are interested in how Vision-Language Models (VLMs) behave in the presence of semantic violations in visual scenes when referring to objects. We focus on object naming, that is, the task of determining the category of an object, a fundamental part of Referring Expression Generation (REG; Krahmer and Van Deemter, 2012). The literature on scene perception suggests that scene semantics mediates human viewers’ expectations about the objects in a scene, and viewers deploy their attentional resources accordingly. Scene semantics also influence how humans describe objects in scenes Hwang et al. (2011); de Groot et al. (2017, 2015). Do VLMs similarly utilise scene-level visual information when generating descriptions of objects? To study this, we compare the naming of objects in semantically (in)congruent scenes. If scene-level information guides model choices, then there should be systematic differences between these two conditions in the way models refer to target referents. While previous work has shown that visual context has a facilitating effect in REG (Junker and Zarrieß, 2024), it does not control for the impact of semantic (in)congruity, and relies on small-scale, purposely-trained models or on artificial images (Junker and ZarrieSS, 2025). To address these issues, we develop a new dataset, Common Objects Out-of-Context (COOCo), which manipulates natural images to introduce objects of varying degrees of consistency with the scene (Figure 1). Similarly to Junker and Zarrieß (2024), we further manipulate the visibility of the target referent and the context by introducing different degrees of noise. Furthermore, since different objects can be related to a scene to different degrees, we model object-scene relatedness as a graded phenomenon. We summarise the contributions of our study as follows.
1. We present COOCo, a novel dataset to evaluate multimodal models’ ability to integrate object-level and scene-level visual information, under both congruent and incongruent conditions.
2. We investigate the impact of scene context on object identification on several state-of-the-art VLMs. We show that under conditions where target objects violate scene semantics, scene context acts as a distractor, while it has a clear facilitating effect when the target is congruent with the scene. This also makes models resilient to noise when it is applied to the target region.
3. Having established that object-scene relatedness exerts an effect across models, we then focus on a single, representative VLM. Using COOCo and another, recently introduced dataset, we investigate how it deploys attention across scene components under different conditions.
@@
Figure 1: An original image of an office scene from the $CO$ dataset and a modified version from COOCo, replacing the laptop with a semantically inconsistent object. Bottom: versions of the modified image with Gaussian noise mask (level = 0.5) on the context and on the target.
@@ 2 Related Work
Context in Referring Expression Generation
@@
Referring Expression Generation (REG) is the task of generating identifying descriptions for a target object. Early REG approaches relying on handcrafted algorithms Krahmer and Van Deemter (2012) incorporated contextual influences by considering the properties that distinguish a target referent from its distractors, with some research also incorporating constraints to guide coherent categorisation for references to plurals, and the use of landmarks for spatial expressions Schüz et al. (2023). More recent neural models for visual REG learn mappings from raw perceptual inputs to linguistic expressions. Here too there has been limited attention on the ability of models to leverage global scene information Schüz et al. (2023). This gap has motivated recent investigations into how scene information may enhance robustness in challenging settings Schüz and ZarrieSS (2023); Junker and ZarrieSS (2025). In particular, Junker and Zarrieß (2024) show that models rely on scene context to generate references even when the target object is fully occluded, indicating that scene context serves both a supportive and disambiguating role during reference generation. Junker and ZarrieSS (2025) further find that scene context influences how VLMs describe ambiguous tangram shapes, though to a lesser extent than humans.
Context in Human Object Recognition
@@
Human object recognition is strongly influenced by contextual expectations Bar (2004); Oliva and Torralba (2007); Võ (2021); Peelen et al. (2023), including both global scene properties (e.g., gist, layout) and local cues (e.g., anchor objects, co-occurring items) Lauer et al. (2018, 2021); Boettcher et al. (2018). Evidence that human overt attention in visual search is semantically constrained comes from viewing tasks, where viewers fixate more on semantically informative (but not necessarily visually salient) regions, and semantic similarity is a strong predictor of gaze trajectories Henderson et al. (2019); Hayes and Henderson (2021). Both highly expected and highly incongruent objects attract more attention than moderately fitting ones Damiano et al. (2024). Motivated by this work, we take a graded view of object-scene semantic similarity in the design of our dataset.
Datasets
@@
Widely used vision and language datasets for referring expression and image caption generation, such as COCO Lin et al. (2015), RefCOCO Kazemzadeh et al. (2014); Yu et al. (2016) and VisualGenome Krishna et al. (2016), lack annotations for object-scene relatedness. Scene violation datasets created within the vision community include ObjAct Shir et al. (2021), Event Task Dresang et al. (2019), Out-of-Context Bomatter et al. (2021), Cut-and-paste dataset Yun et al. (2021), and SCEGRAM Öhlschläger and Võ (2016). However, these datasets tend to be relatively small-scale and do not incorporate degrees of semantic violation or occlusion. In this work, we present a new, large-scale dataset for semantic violations in visual scenes, building on existing resources such as COCO-Search18 Chen et al. (2021), THINGSPlus Hebart et al. (2019); Stoinski et al. (2023) and SUN Xiao et al. (2010). Most related to our work are two recent and concurrently released datasets: VISIONS (Allegretti et al., 2025, further described in § 3) incorporates semantic and spatial violations but remains small in scale (1,136 images), providing a single incongruent manipulation per scene. COinCO Yang et al. (2025) introduces one inpainted object replacement per image and adopts a binary notion of contextual fit based on judgements generated with Molmo Deitke et al. (2024). In contrast, COOCo offers multiple controlled variants of each scene by systematically modulating object–scene relatedness on a continuous scale, enabling finer semantic manipulation than prior datasets. We also evaluate on VISIONS because it is fully human-validated, provides a continuous object-scene relatedness measure, and includes modal object names.
@@ 3 Dataset Creation
We introduce the Common Objects Out-of-Context (COOCo) dataset. The dataset provides a collection of images where the semantic coherence of a scene is modified by the presence of a target object with low, medium or high semantic relatedness to the scene type. COOCo builds upon a framework previously established in the visual perception literature with the SCEGRAM dataset Öhlschläger and Võ (2017), scaling it up to meet the needs of NLP and computer vision research through a novel methodology.
@@
We begin with the publicly available subset of COCO-Search18 Chen et al. (2021), which includes bounding boxes for target objects. This dataset excludes images with people or animals, categories which are known to significantly influence human visual attention  Clarke et al. (2013); End and Gamer (2017). Excluding such cues, therefore, avoids confounding factors and enables us to study object-scene semantic relationships. COCO-Search18 further excludes images with targets with multiple object instances, controlling for target size, location, and image aspect ratio. We retained the 2,241 images where the target object is present.
@@
For each image, we predict the scene label using a Vision Transformer (ViT) [^1] https://huggingface.co/google/vit-base-patch16-224 Dosovitskiy et al. (2021) finetuned on SUN-397 Xiao et al. (2010). We then reduce the SUN label set (397 labels) to 29 (listed in Appendix A.1), selected to balance frequency and semantic diversity and reclassified the images with the same model, excluding all but the selected labels. This approach mitigates risks of category under-representation and excessive variance. This step provided a scene label for each image.
@@
We developed an inpainting pipeline to create versions of images where the target object is replaced with alternatives of varying congruence with the scene (low and medium). We also retain the original image, and additionally include controls where (a) the target is replaced with an object that has high relatedness to the scene, and (b) the target is replaced with a new object of the same category as the original. The latter are included to verify if the effects found in the experiments are due to artefacts caused by inpainting in general, or due to the introduction of incongruent object types. Figure 2 shows an example of a scene and its variants in COOCo.
New objects are determined using human-generated norms in THINGSplus Stoinski et al. (2023), which includes typicality ratings (the degree of representativeness of an object for higher-level categories) and ratings of the real-world size of objects. We retain only typical objects (typicality ranging from 0.3 to 1) and remove animate objects to ensure consistency with COCO-Search18. We then match COCO-Search18 targets to candidate objects in THINGSPlus with similar real-world size norms, excluding any candidates differing in size from the targets by more than 25 units.
We compute semantic relatedness between candidate replacements and scene labels following Hayes and Henderson (2021), who found strong correlations between embeddings in ConceptNet Numberbatch Speer et al. (2018) and human visual attention shifts. We use the cosine similarity between scene and object label embeddings to filter low-, medium- and high-similarity candidates.
@@
Given an image, a scene label and a target object with its bounding region and segmentation-based mask obtained from COCO-Search18, we remove the target from the image, and generate 15 new versions of the scene where the target is replaced with objects having low, medium or high similarity to the scene label (see Figure 2). We also generate an image containing the target as in the original image as a control condition. Only images that pass a verification test are retained. Full details of the inpainting pipeline and verification tests are in Appendix A.1.
Final COOCo dataset
The final COOCo dataset comprises 18,395 images: 1,862 original images with corresponding object-removed versions; 5,480 generated images with low relatedness; 5,572 with medium relatedness; 1,771 with high relatedness; and 1,848 images for the same-target condition.
VISIONS dataset
We conduct a subset of our experiments on the VISIONS dataset Allegretti et al. (2025) and compare the results obtained on COOCo with those on it. VISIONS comprises 1,136 naturalistic indoor images constructed from 165 scene setups across seven scene categories: bathroom, bedroom, church, kitchen, office, restaurant, and waiting room. Each setup is photographed in four principal versions created by manipulating the semantic consistency (consistent vs. inconsistent) of a target object, identified with bounding box coordinates, and its spatial position (left vs. right). [^2] Although the dataset also includes additional swap-based variants for 119 scenes where the target object was exchanged with another object in the scene to produce four mirrored counterparts of the canonical versions, we use only these four canonical versions, as swap manipulations are not part of our analyses. VISIONS provides an extensive set of perceptual and conceptual norms for both objects and scenes, including object-consistency ratings, which quantify how contextually appropriate the critical object is in each scene. We utilise these ratings. VISIONS also provides human-annotated target objects with free-text labels; we use the modal name for each object as the reference label in our analysis, together with the original object label defined by the dataset creators.
@@ 4 Experimental Setup
We design our experiments to address how VLMs leverage contextual information for object naming, an essential component of Referring Expression Generation. (1) First, we evaluate to what extent VLMs rely on scene context under conditions of varying degrees of semantic congruence between an object and its background; and on how robustly they perform REG under different levels and configurations of visual noise. These experiments are conducted on a variety of SOTA VLMs on the COOCo dataset. Each model is provided with an image and a specified target region, and is prompted to name the object within that area using a structured text prompt. Anticipating the outcomes of this analysis (see § 5), we identify a number of trends across models, showing that they are sensitive to scene semantics when referring to targets and that scene context supports REG when target objects are obscured by noise (cf. Junker and Zarrieß, 2024). (2) Having established these trends across multiple models, we select a single performant model for in-depth analysis of its outputs (§ 6.1) and its attention deployment across different experimental settings (§ 6.2). This analysis is performed on our own dataset, COOCo, with replications on the VISIONS dataset Allegretti et al. (2025). This allows us both to generalise findings and control for possible artefacts introduced by inpainting in COOCo.
@@ 4.1 Models
Molmo Deitke et al. (2024) (7B) is trained on a dataset including ca. 2.3M human-annotated pairs of questions and references to image regions based on 2D points, from 428k images.
@@
LLaVA-OneVision Li et al. (2025) (0.5B and 7B) was trained on data with bounding boxes, incorporating RefCOCO Kazemzadeh et al. (2014); Yu et al. (2016)[>50k samples] and Visual Genome Krishna et al. (2016)[86k] for single-image settings, and multi-image datasets including WebQA Chang et al. (2022)[9.3k]. The model is instruction-tuned to accept spatial information directly in text input in the form of normalised bounding box coordinates.
@@
For detailed descriptions of preprocessing steps and region-of-interest (ROI) prompt formats for each model, refer to Appendix A.2.
Prompts
We test models with structured prompts (details in Appendix A.2) with an explicit reference to a region, for example: What is the object in this part of the image $x_{1}$, $y_{1}$, $x_{2}$, $y_{2}$]? Answer with the object’s name only. No extra text.
4.2 Noise injection
We add Gaussian noise ($\lambda=0.0,0.5,1.0$), ranging from mild distortion to maximum information loss, to the target region, the context region or the entire image (‘All noise’). Pixel values in the affected area are perturbed with noise sampled from a normal distribution and clipped to stay within valid image ranges.
@@ 4.3 Metrics
RefCLIPScore Hessel et al. (2022) is the harmonic mean between (1) the CLIPScore Hessel et al. (2022), which captures the semantic alignment between the generated caption and the image; and (2) the highest cosine similarity between the CLIP Radford et al. (2021) embeddings of the generated caption and any reference caption. This quantifies the alignment between a generated description for a target and both its visual features (extracted from its image patch) and correct label.
Text-Based Semantic Similarity is computed as the cosine similarity between CLIP text embeddings. We use this measure to assess the semantic similarity between the model’s output and the target label (from which Accuracy is derived), as well as between the model’s output and the scene label in § 6.1.
Accuracy, defined as 1 if the cosine similarity between the model’s output and the target label exceeds a threshold (0.9), and 0 otherwise. To illustrate, consider the target car. While the output car scores 1, related terms bus (0.86), truck (0.89), and police car (0.89) score 0. In contrast, variants such as SUV (0.90), a race car (0.90), and the part–whole expression car door (0.90) surpass 0.9 and are counted as correct. [^3] We also computed a hard accuracy metric based on Levenshtein similarity, which quantifies the character-level edit distance between generated and reference captions. Since results are very similar to those obtained with the cosine-based metric, we do not report them here.
5 Results across models
In this section, we report and discuss the main results across all models and experimental conditions in COOCo.
Figure 3: RefCLIPScores per model at noise level 0 across relatedness conditions. Models perform best in the original and same target conditions, with performance declining as target relatedness decreases.